OpenStudy (anonymous):

ah help:( A home furnishings store sells cylindrical storage containers in various sizes. The smallest is 5 inches tall and has a radius of 2 inches. Find the volume of the container described.

OpenStudy (diyadiya):

\[Volume \ of \ Cylinder = \pi \times (radius)^2 \times height \]

OpenStudy (anonymous):

OH. hahaha okay thankyou!

OpenStudy (anonymous):

20π